# Haoyang Chen1,2,3  Dongfang Sun1,2111These authors contributed equally to this work.  Caoyuan Ma1,2111These authors contributed equally to this work.  Shiqin Wang1,2  Kewei Zhang1,2

### v2

@@ 4.2 Comparisons
We conducted a comprehensive quantitative and qualitative evaluation of the proposed method and competing methods on the CMC and FMC datasets. Our competing methods include the training-based approaches ControlNet [33], T2I-Adapter [19], and Uni-ControlNet [37], as well as the training-free approaches Multi-Sketch [4], PNP [27], and FreeControl [17]. The quantitative results are presented in Table 1 and qualitative results are shown in Figure 5.
We observe that training-based methods [33, 19, 37] are highly susceptible to subjective biases in sketches and often fail to adhere to text prompts, producing distorted and unrealistic images with lower CLIP-T scores. Training-free methods [4, 8, 17] suffer from concept omissions, conspicuous artifacts, and inconsistencies both among concepts and between concepts and their backgrounds, resulting in poor image realism. These issues stem from mutual interference during the “snapshot” generation of multiple concepts. In contrast, our proposed method faithfully reconstructs real-world scenes and achieves the best performance across all evaluation metrics on both the CMC and FMC datasets, demonstrating its effectiveness.
@@ 4.4 User Preference Experiments
We invited 12 participants (group A) to provide subjective readouts and an additional 28 participants (group B) to evaluate the reconstructed images. Participants in group A were first asked to select a deeply impressive scene from their photo gallery and then provide a textual description along with sketches within two minutes. Our method and competing approaches, including two baselines (text-reward optimization and text-reward optimization with sketch-guided optimization based on all concepts) as well as an optimal method, FreeControl [17], were used to generate images from these readouts. Participants in both groups were then shown the subjective readouts, the reference image, and the generated images for the same scene. They were asked to answer four single-choice questions in a single-blind manner: Q1: Which image is closer to the original image (both style and content)? Q2: Which image has higher visual quality (e.g., more realistic, more natural)? Q3: Which image better matches the structure of the sketch? Q4: Which image better matches the textual description?
Figure 7 presents the statistics of the collected results. Participants judged our reconstructions as being closer to the corresponding real-world scenes compared to competing methods. Similar conclusions were observed for the other three questions. However, there was an exception in the comparison between our method and text-reward optimization regarding image quality (Q2). We observed that text-reward optimization without additional control tends to hallucinate details that enhance perceived realism, leading some users to prefer it for Q2. This also explains why our advantage over text-reward optimization in Q1 was not statistically significant and why some participants expressed hesitation in this case.
@@ 5 Conclusions
This paper presented Subjective Camera, a paradigm for reconstructing real-world scenes from humans’ mental impressions. Our framework relies on the most accessible readouts of mental imagery, i.e., textual descriptions and sketches, and employs a training-free, sequence-aware optimization process. This approach eliminates the need for costly fine-tuning of text-to-image diffusion models on large-scale data while mitigating challenges such as dependence on sketch quality and geometric misalignment. Extensive experiments on both synthetic and real-world sketch datasets demonstrate that Subjective Camera achieves state-of-the-art performance, outperforming both training-free and training-based sketch-to-image generation methods across all evaluation metrics. Furthermore, user preference studies confirm that participants consistently favor our method.
In closing, we address a natural question: “With the rapid advancement of generative AI, will physical cameras ever become unnecessary” For now, the answer is no. First, the reconstructed images still fall short of perfectly matching actual scenes. This limitation largely stems from the fact that current T2I models are trained on highly diverse data, creating gaps between the prior and the specific scene to be reconstructed. One promising direction to mitigate this is to personalize the prior models using user-specific galleries. Second, we observe that variations in the order and quality of sketches have a significant impact on reconstruction results. To address this, we plan to enhance our framework with computational tools that guide users in providing more precise and suitable subjective readouts. For these reasons, we designate this system as Subjective Camera 0.1, leaving these improvements for future iterations.
Acknowledgements This work was supported in part by the National Natural Science Foundation of China (62225113, 624B2109, 623B2079) and the Zhongguancun Academy Project (20240308). We thank Po-Hsun Huang for providing the questionnaire system for our user study, Kaijin Zheng for assistance with data processing, and all participants who took part in the studies.

### v3

@@ 4.2 Comparisons
We conducted a comprehensive quantitative and qualitative evaluation of the proposed method and competing methods on the CMC and FMC datasets. Our competing methods include the training‑based approaches ControlNet [33], T2I‑Adapter [19], and Uni‑ControlNet [37], as well as the training‑free approaches Multi‑Sketch [4], PNP [27], and FreeControl [17]. The quantitative results are presented in Table 1 and qualitative results are shown in Figure 5.
We observe that training‑based methods [33, 19, 37] are highly susceptible to subjective biases in sketches and often fail to adhere to text prompts, producing distorted and unrealistic images with lower CLIP-T scores. Training‑free methods [4, 8, 17] suffer from concept omissions, conspicuous artifacts, and inconsistencies both among concepts and between concepts and their backgrounds, resulting in poor image realism. These issues stem from mutual interference during the “snapshot” generation of multiple concepts. In contrast, our proposed method faithfully reconstructs real‑world scenes and achieves the best performance across all evaluation metrics on both the CMC and FMC datasets, demonstrating its effectiveness.
@@ 4.4 User Preference Experiments
We invited 12 participants (group A) to provide subjective readouts and an additional 28 participants (group B) to evaluate the reconstructed images. Participants in group A were first asked to select a deeply impressive scene from their photo gallery and then provide a textual description along with sketches within two minutes. Our method and competing approaches, including two baselines (text‑reward optimization and text‑reward optimization with sketch‑guided optimization based on all concepts) as well as an optimal method, FreeControl [17], were used to generate images from these readouts. Participants in both groups were then shown the subjective readouts, the reference image, and the generated images for the same scene. They were asked to answer four single‑choice questions in a single‑blind manner: Q1: Which image is closer to the original image (both style and content)? Q2: Which image has higher visual quality (e.g., more realistic, more natural)? Q3: Which image better matches the structure of the sketch? Q4: Which image better matches the textual description?
Figure 7 presents the statistics of the collected results. Participants judged our reconstructions as being closer to the corresponding real‑world scenes compared to competing methods. Similar conclusions were observed for the other three questions. However, there was an exception in the comparison between our method and text‑reward optimization regarding image quality (Q2). We observed that text‑reward optimization without additional control tends to hallucinate details that enhance perceived realism, leading some users to prefer it for Q2. This also explains why our advantage over text‑reward optimization in Q1 was not statistically significant and why some participants expressed hesitation in this case.
@@ 5 Conclusions
This paper presented Subjective Camera, a paradigm for reconstructing real‑world scenes from humans’ mental impressions. Our framework relies on the most accessible readouts of mental imagery, i.e., textual descriptions and sketches, and employs a training‑free, sequence‑aware optimization process. This approach eliminates the need for costly fine‑tuning of text‑to‑image diffusion models on large‑scale data while mitigating challenges such as dependence on sketch quality and geometric misalignment. Extensive experiments on both synthetic and real‑world sketch datasets demonstrate that Subjective Camera achieves state‑of‑the‑art performance, outperforming both training‑free and training‑based sketch‑to‑image generation methods across all evaluation metrics. Furthermore, user preference studies confirm that participants consistently favor our method.
In closing, we address a natural question: “With the rapid advancement of generative AI, will physical cameras ever become unnecessary” For now, the answer is no. First, the reconstructed images still fall short of perfectly matching actual scenes. This limitation largely stems from the fact that current T2I models are trained on highly diverse data, creating gaps between the prior and the specific scene to be reconstructed. One promising direction to mitigate this is to personalize the prior models using user‑specific galleries. Second, we observe that variations in the order and quality of sketches have a significant impact on reconstruction results. To address this, we plan to enhance our framework with computational tools that guide users in providing more precise and suitable subjective readouts. For these reasons, we designate this system as Subjective Camera 1.0, leaving these improvements for future iterations.
Acknowledgements This work was supported in part by the National Natural Science Foundation of China (62225113, 624B2109, 623B2079) and the Zhongguancun Academy Project (20240308). We thank Po‑Hsun Huang for providing the questionnaire system for our user study, Kaijin Zheng for assistance with data processing, and all participants who took part in the studies.